OpenStudy (anonymous):

need help solving: 8x^2=9x using quadric formula

OpenStudy (anonymous):

\[\frac{-(-9) \pm \sqrt{(-9)^2 - 4(8)(0)}}{2(8)}\]

OpenStudy (chaise):

Once again my equation editor isn't working - sorry. Make the equation equal to zero: 8x^2=9x 8x^2-9x + 0 = 0 ({-b +/- sqrt{b^2-4ac)/{2a} a = 8 b = -9 c = 0 Substitute your values in: x = 9/8 x = 0

OpenStudy (chaise):

Maths boy has kindly substituted in your values :)

OpenStudy (anonymous):